### v3

# Addressing divergent representations from causal interventions on neural networks

Abstract
A common approach to mechanistic interpretability is to causally manipulate model representations via targeted interventions in order to understand what those representations encode. Here we ask whether such interventions create out-of-distribution (divergent) representations, and whether this raises concerns about how faithful their resulting explanations are to the target model in its natural state. First, we demonstrate theoretically and empirically that common causal intervention techniques often do shift internal representations away from the natural distribution of the target model. Then, we provide a theoretical analysis of two classes of such divergences: ‘harmless’ divergences that occur in the null-space of the weights and from covariance within behavioral decision boundaries, and ‘pernicious’ divergences that activate hidden network pathways and cause dormant behavioral changes. Finally, in an effort to mitigate the pernicious cases, we apply and modify the Counterfactual Latent (CL) loss from Grant (2025) allowing representations from causal interventions to remain closer to the natural distribution, reducing the likelihood of harmful divergences while preserving the interpretive power of the interventions. Together, these results highlight a path towards more reliable interpretability methods.
 [^1] Link to public repository: https://github.com/grantsrb/rep_divergence
1 Introduction
A central goal of mechanistic interpretability is to understand what the internal representations of neural networks (NNs) encode and how this gives rise to their behavior. Perhaps the most powerful approach to pursuing this goal is through causal interventions, where methods such as activation patching and Distributed Alignment Search (DAS) directly manipulate internal representations to test how they affect outputs (geiger2021; geiger2023boundless; wu2024alpacadas; wang2022patching; meng2023activpatching; nanda2022attrpatching; csordás2024). Indeed, even correlational methods such as Sparse Autoencoders (SAEs) and Principal Component Analysis (PCA) often use causal interventions as a final judge for whether the features they identify are truly meaningful (huang2024ravel; dai2024representational). Causal interventions thus occupy a central place in making functional claims about neural circuitry (pearl2010causalmediation; geiger2024causalabstractiontheoreticalfoundation; geiger2025causalunderpin; lampinen2025representation; braun2025functionalvsrepsim).
Figure 1: Causal interventions can recruit hidden circuits that produce misleadingly confirmatory or dormant behavior. (a) Consider natural pathways (dashed arrows) for two classes A and B that carry activity to different behavioral outputs $y$. In a hypothetical intervention meant to find path A, patching $h^{1}$ with a divergent representation can activate distinct, hidden pathways (solid arrows) that result in misleadingly confirmatory behavior (orange) and/or undetected behavior (red). (b) Consider 2D projections of the neural activity of $h^{1}$ for a different network that classifies states into one of 10 classes (denoted by hue). Suppose that natural representations (dark points) lie within well-defined decision boundaries (dashed lines) and covary along causal axes, and that intervened representations (light points) are constructed by patching the first axis from a sampled natural representation. Although these representations diverge from the natural distribution, this can be harmless (top) or pernicious (bottom) depending on the network’s functional landscape. In particular, it can be pernicious if the network has a functional landscape where intervened activity unknowingly recruits hidden circuits (visualized as an orange region) or crosses dormant behavioral boundaries (red regions).
The use of causal interventions rests on a fundamental assumption that counterfactual model states created by interventions are realistic for the target model. Despite its pervasiveness, however, this assumption is often untested. For example, some activation patching experiments multiply feature values by up to 15x (lindsey2025biology); in these settings, it seems possible that intervened representations diverge significantly from the NN’s natural distributions. This raises questions about the reliability of causal interventions for mechanistic interpretability. Do divergent representations change what an intervention can say about an NN’s natural mechanisms? When, and to what extent, is it okay for such divergences to occur? When it is not okay, how can we prevent them from occurring?
In this work, we provide both empirical and theoretical insight on these issues. We first demonstrate that divergent representations are a common issue for causal interventions – across a wide range of intervention methods, we find that intervened representations often do diverge from the target NN’s natural distribution. We then provide theoretical examples of two types of divergence: ‘harmless’ divergences that can occur from within-decision-boundary covariance along causal dimensions or from deviations in the null-spaces of the NN layers, and ‘pernicious’ divergences that activate hidden network pathways and can cause dormant changes to behavior. We provide discussion on how harmless and pernicious cases are not always mutually exclusive, where the harm often depends on the specific mechanistic claims. Finally, we provide a broad-stroke, initial solution for mitigating pernicious divergences by minimizing all intervened divergences. We show that we can use the Counterfactual Latent (CL) auxiliary loss introduced in grant2025mas to reduce all representational divergence in the Boundless DAS setting from wu2024alpacadas while maintaining the same behavioral accuracy; and we introduce a modified version of the CL loss that targets causal subspaces and show that it can improve out-of-distribution (OOD) intervention performance on synthetic tasks. Although we do not propose this method as the final solution to representational divergence, we pose it as a step towards more reliable interventions.
We summarize our contributions as follows:
  1. 1.
    We show theoretical and empirical examples of divergence between natural and causally intervened representations for a variety of causal methods (Section 3).
  2. 2.
    We provide a theoretical treatment of cases in which divergence can arise innocuously from variation in the null-space and from covariance within decision boundaries of behavioral subspaces, demonstrating that some divergences can be okay and even desired (Section 4.1)
  3. 3.
    We provide synthetic examples of cases where divergent representations can (1) activate hidden computational pathways while still resulting in hypothesis-affirming behavior, and (2) cause dormant behavioral changes, together raising questions about the mechanistic claims that can be made from patching results alone (Section 4.2).
  4. 4.
    We use the CL auxiliary loss from grant2025mas to minimize patching divergence directly in the Large Language Model (LLM) Boundless DAS experimental setting from wu2024alpacadas, and we introduce a stand-alone, modified CL loss that exclusively minimizes divergence along causal dimensions, improving OOD intervention performance in synthetic settings. Together, these results provide an initial step towards preventing pernicious divergences (Section 5).
2 Background and Related Work
2.1 Activation Patching
Activation patching generally refers to a process of “patching” (i.e. substituting) some portion of neural activity at an intermediate layer into or from a corrupted forward pass of a network (geiger2020neural; vig2020causal; wang2022patching; meng2023activpatching; zhang2024bestpracticepatching). It can be performed at various granularities such as whole layers, attention heads, or individual neurons. Many forms of activation patching can be unified under the assumption that subspaces, rather than individual neurons, are the atomic units of NN representations (PDP1; PDP2; Smolensky1988; elhage2022superposition; geiger2021; grant2025das). Activation patching at the level of individual neurons can be understood as subspace patching along neuronal axes, and many of its high-level granularities can be understood as specific forms of individual neuron patching (geiger2020neural; vig2020causal; wang2022patching; meng2023activpatching).
2.2 Distributed Alignment Search
Distributed Alignment Search (DAS) (geiger2021; geiger2023boundless; wu2024alpacadas) can be understood as a form of activation patching that operates in a transformed basis so that specific, causally relevant subspaces can be manipulated analogously to high-level variables from causal abstractions (CAs) (e.g. symbolic programs). Many cases of individual neuron (coordinate) patching can be understood as specific cases of DAS that use the identity transform. We use DAS in Sections 4 and 5, so here we introduce its theory and background.
DAS finds alignments between neural representational subspaces and causal variables from Causal Abstractions (CAs) by testing the hypothesis that an NN’s latent state vector $h\in R^{d_{m}}$ can be transformed into a vector $z\in R^{d_{m}}$ that consists of orthogonal subspaces encoding interpretable variables. This transformation is performed by a learnable, invertible Alignment Function (AF), $z=\mathcal{A}(h)$. We restrict our considerations to linear AFs of the form $\mathcal{A}(h)=Wh$ where $W\in R^{d_{m}\times d_{m}}$ is invertible. This transformation allows us to formulate $h$ in terms of interpretable variables and to manipulate encoded values.
For a given CA with variables $\text{var}_{i}\in\{\text{var}_{1},\text{var}_{2},...,\text{var}_{n}\}$, DAS tests the hypothesis that $z$ is composed of subspaces $\vec{z}_{\text{var}_{i}}\in R^{d_{\text{var}_{i}}}$ corresponding to each of the variables from the CA. A causally irrelevant subspace $\vec{z}_{\text{extra}}\in R^{d_{\text{extra}}}$ is also included to encode extraneous, functionally irrelevant neural activity (i.e., the behavioral null-space).
$\mathcal{A}(h)=z=\begin{bmatrix}\vec{z}_{\text{var}_{1}}\\ \cdots\\ \vec{z}_{\text{var}_{n}}\\ \vec{z}_{\text{extra}}\end{bmatrix}$ (1)
Here, each $\vec{z}_{\text{var}_{i}}\in R^{d_{\text{var}_{i}}}$ is a column vector of potentially different lengths, where $d_{\text{var}_{i}}$ is the subspace size of $\text{var}_{i}$, and all subspace sizes satisfy $d_{\text{extra}}+\sum_{i=1}^{n}d_{\text{var}_{i}}=d_{m}$. The value of a single causal variable encoded in $h$ can be manipulated through an interchange intervention defined as follows:
$\hat{h}=\mathcal{A}^{-1}((\mathcal{I}-D_{\text{var}_{i}})\mathcal{A}(h^{\text{trg}})+D_{\text{var}_{i}}\mathcal{A}(h^{\text{src}}))$ (2)
Here, $D_{\text{var}}\in R^{d_{m}\times d_{m}}$ is a manually defined, block diagonal, binary matrix that defines the subspace size $d_{\text{var}_{i}}$, and $\mathcal{I}\in\mathbb{R}^{d_{m}\times d_{m}}$ is the identity matrix. Each $D_{\text{var}_{i}}$ has a set of $d_{\text{var}_{i}}$ contiguous ones along its diagonal to isolate the dimensions that make up $\vec{z}_{\text{var}_{i}}$. $h^{\text{src}}$ is the source vector from which the subspace activity is harvested, $h^{\text{trg}}$ is the target vector into which the harvested activity is patched, and $\hat{h}$ is the resulting intervened vector that replaces $h^{\text{trg}}$ in the model’s processing. This allows the model to make predictions using a new value of variable $\text{var}_{i}$.
To train $\mathcal{A}$, DAS uses counterfactual behavior $c\sim\mathcal{D}$ as training labels, where $c$ is generated from the CA. $c$, for a given state of a CA and its context, is the behavior that would have occurred had a causal variable taken a different value and everything else remained the same. $c$ is generated by freezing the state of the environment, changing one or more variable values in the CA, and using the CA to generate new behavior in the same environment using the new values. We train $\mathcal{A}$ on intervention samples while keeping the model parameters frozen, minimizing the following objective (for non-sequence-based settings):
$\mathcal{L}_{\text{DAS}}(\mathcal{A})=-\frac{1}{N}\sum_{k=1}^{N}\log p_{\mathcal{A}}\!\left(c^{(k)}\,\middle|\,x^{(k)},\hat{h}^{(k)}\right),$ (3)
where $N$ is the number of samples in the dataset, $c^{(k)}$ is the counterfactual label in sample $k$, $x^{(k)}$ is the model input data, and $p_{\mathcal{A}}(\cdot\mid\cdot)$ is the model’s conditional probability distribution given the intervened latent vector, $\hat{h}$. We minimize $\mathcal{L}_{\text{DAS}}(\mathcal{A})$ using gradient descent, backpropagating into $\mathcal{A}$ with all model weights frozen. $\mathcal{A}$ is evaluated on new intervention data, where the model’s accuracy on $c$ following each intervention is referred to as the Interchange Intervention Accuracy (IIA).
2.3 Problematic Causal Interventions
Prior work has implicitly explored issues related to representational divergence from causal interventions. For example, methods such as causal scrubbing or noising/denoising activation patching (wang2022patching; lawrencec2022causalscrubbing; meng2023activpatching; chen2025completenessOrAndAdder; zhang2024bestpracticepatching) intentionally introduce divergent representations to test the sufficiency, completeness, and faithfulness of proposed circuits. Works such as wattenberg2024featuremultiplicityanddarkmatterinNNs, meloux2024causalabstractionsnotunique, and chen2025completenessOrAndAdder also implicitly explore the dangers of divergent intervened representations by showing how circuits and features can be redundant or have combinatorial effects that are difficult to enumerate given current methodologies, while zhang2024bestpracticepatching and heimersheim2024practicepatching point out easy misinterpretations of patching results. shi2024circuithypothesistesting and wang2022patching provide criteria centered on faithfulness, completeness, and minimality for evaluating circuits through causal interventions. A body of work on counterfactual explanations exists, some of which has explored differences between on-manifold and off-manifold adversarial attacks (stutzheinschiele2019adversarialmanifolds), and some works have explored constraining counterfactual features to the manifold of the dataset (verma2024counterfactualexplanationreview; tsiourvas2024manifoldalignedcounterfactuals). Our proposed method in Section 5 differs in that it trains a principled alignment to generate counterfactual representations and it constrains deviations along causal dimensions. For DAS in particular, makelov2023interpillusions demonstrate that it is possible to produce an interaction between the null-space and dormant subspaces that affect behavior. Because they define dormant subspaces as those that do not vary across different model inputs, variation along these directions is, by definition, a form of divergent representation. Finally, sutter2025nonlinrepdilemma posit that it is possible to align any causal abstraction to NNs under a number of assumptions including a sufficiently powered, non-linear alignment function, raising questions about what non-linear causal interventions really tell us.
3 Are divergent representations a common phenomenon?
We begin by demonstrating that divergent representations are a common (if not likely) outcome of causal interventions, both in theory and in practice. We do not yet consider its perniciousness, however, and we reserve the question of whether and when divergence is harmful for Section 4.
3.1 For most manifolds, coordinate patching guarantees divergence
We first consider a theoretical setting where coordinate-based patching of one or more vector dimensions is performed on a single manifold, similar to what might be done in neuron level activation patching (vig2020causal; geiger2021). We prove that in this setting, divergent representations are guaranteed to occur if patching is performed exhaustively. For simplicity, we consider a minimal version of this proof that involves a circular manifold with two dimensions. A more general proof, which applies to most manifold geometries, can be found in Appendix A.2.
Figure 2: Representational divergence is a common occurrence across various interventions. (a) Directly replacing a coordinate value in one natural representation (orange) with the value from another will eventually create divergent representations (blue). (b) Top two principal components of natural and corresponding intervened representations, taken from the residual stream at the intervention position and with PCA is performed over the combined set of natural and intervened vectors, for three popular causal intervention techniques: a replication of feng2024languagemodelsbindentities for mean difference patching, reconstructed vectors for a single transformer layer using SAELens (bloom2024saelens) for sparse autoencoder, and interchange interventions for Boundless DAS (wu2024alpacadas). (c) L2 distance between natural and corresponding intervened representations, and Earth Mover’s Distance (EMD) between natural and intervened distributions (with baseline comparing the natural distribution to itself).
Formally, let $\mathcal{M}_{K}=\{\,c_{K}+u:\|u\|_{2}\leq r_{K}\,\}\subset\mathbb{R}^{2}$ be a class-$K$ manifold with centroid $c_{K}\in\mathbb{R}^{2}$ and radius $r_{K}>0$. Given two native representations $h^{\textit{src}}=c_{K}+u$ and $h^{\textit{trg}}=c_{K}+v$, let us define a coordinate patch (onto class $K$) that keeps the first coordinate of $h^{\textit{src}}$ and the second of $h^{\textit{trg}}$:
$\hat{h}\;=\;\begin{bmatrix}h_{1}^{\textit{src}}\\ h_{2}^{\textit{trg}}\end{bmatrix}=\begin{bmatrix}c_{K,1}+u_{1}\\ c_{K,2}+v_{2}\end{bmatrix}$.
Then the deviation from $c_{K}$ is
$\hat{h}-c_{K}\;=\;(u_{1},\,v_{2})^{\top},\qquad\|\hat{h}-c_{K}\|_{2}^{2}\;=\;u_{1}^{2}+v_{2}^{2}.$
Proposition (coordinate patching exceeds the class radius). If $h^{\textit{src}},h^{\textit{trg}}\in\mathcal{M}_{K}$ (i.e., $\|u\|_{2}\leq r_{K}$ and $\|v\|_{2}\leq r_{K}$), then the patched point $\hat{h}$ is off-manifold whenever $u_{1}^{2}+v_{2}^{2}>r_{K}^{2}$. In particular, there exist boundary points $h^{\textit{src}},h^{\textit{trg}}\in\partial\mathcal{M}_{K}$ with $u=(r_{K},0)$ and $v=(0,r_{K})$ such that $\|\hat{h}-c_{K}\|_{2}=\sqrt{r_{K}^{2}+r_{K}^{2}}=r_{K}\sqrt{2}>r_{K}$.
Proof. Since $\hat{h}-c_{K}=(u_{1},v_{2})$, we have $\|\hat{h}-c_{K}\|_{2}^{2}=u_{1}^{2}+v_{2}^{2}$. Choosing $u=(r_{K},0)$ and $v=(0,r_{K})$ gives the stated violation. ∎
As noted above, this intuition holds for all manifold shapes other than axis-aligned hyper-rectangles (see Appendix A.2). Thus, in these relatively simple theoretical intervention settings, divergent representations are guaranteed to occur with enough intervention samples.
3.2 Many existing causal methods empirically produce divergence
Stepping beyond the theoretical setting, we also empirically demonstrate that common, real-world causal interpretability methods often produce divergent representations. The notion that these methods can create divergence has been indirectly explored in previous work through attention patterns in Gaussian noise corrupted activation patching (zhang2024bestpracticepatching). Here we extend this view to three popular causal intervention methods: mean difference vector patching (e.g. feng2024languagemodelsbindentities), Sparse Autoencoders (e.g. bloom2024saelens), and DAS (e.g. wu2024alpacadas).
Figure 2 shows the top two principal components of the natural and intervened representations for each intervention method, distinguished by color. These results demonstrate that divergence is a common phenomenon in practice and is not specific to any one method. Even simple methods that patch along a single mean direction are subject to divergence, despite high behavioral accuracy (feng2024languagemodelsbindentities; geiger2023boundless; wu2024alpacadas). We quantify this divergence in Figure 2(c) using the Earth Mover’s Distance (EMD) (villani2009wassersteinEMD) between the full dimensional natural and intervened distributions, using the corresponding natural-natural comparison as a baseline. We see that the intervened divergence exceeds that of the natural (more metrics and details are in Appendix A.1). Note that this result does not necessarily imply that the respective methods are invalid or that their claims are incorrect; the panels are only meant to show the presence of divergence.
4 When are divergent representations harmless or pernicious?
Having demonstrated that divergent representations are a common phenomenon, we now turn our attention to considering whether divergence is a concerning phenomenon. We propose that divergence is harmless to most functional claims if it occurs in the behavioral null-space of the model weights or from covariance within decision boundaries, and that it can be pernicious if it recruits hidden pathways or leads to dormant behavioral changes. However, we stress that the harm is inherently claim dependent, meaning that these divergences are not mutually exclusive.
4.1 Harmless cases
We first explore cases in which we might consider divergent representations to be harmless to many general functional claims. First among these is a set of cases in which divergence is bottlenecked into the null-space of the next computational layer(s). Formally, we define the null-space of a weight matrix $W\in\mathbb{R}^{d_{2}\times d_{1}}$ as $\mathcal{N}(W)=\{v\in\mathbb{R}^{d_{1}}\,\,|\,\,Wv=0\}$. Neural activity in the null-space of the weights refers to any changes $\delta\in\mathbb{R}^{d_{1}}$ for which $W(h^{\ell}+\delta)=Wh^{\ell}$. We can generalize this notion beyond the weights to all remaining computations after and including layer $\ell$ in a neural network denoted $f_{\geq\ell}(h)$. We call this the behavioral null-space for a layer $\ell$, defined as $\mathcal{N}(f_{\geq\ell}(h^{\ell}))=\{v\in\mathbb{R}^{d_{1}}\,\,|\,\,f_{\geq\ell}(h^{\ell}+v)=f_{\geq\ell}(h^{\ell})\}$.
If one cares only about the model’s output behavior, then divergence that is bottlenecked into null-spaces and that therefore has no effect on the model’s behavior would be harmless. Indeed, this covers many present day scenarios, and one could even desire to intervene along null-space directions in order to causally test that such dimensions are indeed in the null-space. It is, however, difficult to define the full behavioral null-space in practice due to the possibility of dormant behavioral changes, which we discuss in Section 4.2.2.
A less obvious case of harmless divergence can occur from covariance within a network’s decision boundaries. For example, the top panel of Figure 1(b) provides a simulated depiction of harmless divergence caused by covariance along causal dimensions. Here, divergent representations are created by swapping the first causal dimension of sampled natural vectors, however, these representations remain within the desired decision boundaries, resulting in the appropriate counterfactual behavior.
We can expand this intuition using a behaviorally binary subspace as an example, where we define a behaviorally binary subspace as a vector subspace which causally impacts the outputs of a future processing layer (e.g., classification labels) only through its sign. Formally, let $f:\mathbb{R}^{d}\to\mathbb{R}^{k}$ denote a computational layer (possibly consisting of multiple NN layers and functions) where $k$ is the output dimensionality. Let $\operatorname{sign}(\cdot)$ denote the elementwise sign map $\operatorname{sign}:\mathbb{R}^{d_{\text{var}}}\to\{-1,1\}^{d_{\text{var}}}$, and assume a fixed alignment function $\mathcal{A}$ and subspace selection matrix $D_{\text{var}}\in\mathbb{R}^{d\times d}$. A linear subspace $Z\subseteq\mathbb{R}^{d}$ is behaviorally binary (with respect to $f$, $D_{\text{var}}$, and $\mathcal{A}$) iff for all $D_{\text{var}}\mathcal{A}(h),D_{\text{var}}\mathcal{A}(h^{\prime})\in Z$,
$\operatorname{sign}(D_{\text{var}}\mathcal{A}(h))=\operatorname{sign}(D_{\text{var}}\mathcal{A}(h^{\prime}))\;\;\Longrightarrow\;\;f(h)=f(h^{\prime})$ (4)
Now, suppose we have an NN with two causal subspaces, $\mathbf{\vec{z}_{\text{var}_{a}}}\subseteq\mathbb{R}^{d_{\text{var}_{a}}}$ and $\mathbf{\vec{z}_{\text{var}_{b}}}\subseteq\mathbb{R}^{d_{\text{var}_{b}}}$, with values $\vec{z}^{\,(x_{i})}_{\text{var}_{a}}$} and $\vec{z}^{\,(x_{i})}_{\text{var}_{b}}$}, for a model input $x_{i}$, where we use the bold, tilde notation to distinguish variables from their values. Furthermore, assume that $\mathbf{\vec{z}_{\text{var}_{b}}}$ is a behaviorally binary subspace that co-varies with, $\mathbf{\vec{z}_{\text{var}_{a}}}$. Using $h^{(x_{i})}$ and $z^{(x_{i})}$ from Equation 1 under a given input $x_{i}$, we use the following definition:
$\mathcal{A}(h^{(x_{i})})=z^{(x_{i})}=\begin{bmatrix}\mathbf{\vec{z}^{\;(x_{i})}_{\text{var}_{a}}}={\vec{z}^{\,(x_{i})}_{\text{var}_{a}}}\\ \mathbf{\vec{z}^{\;(x_{i})}_{\text{var}_{b}}}={\vec{z}^{\,(x_{i})}_{\text{var}_{b}}}\end{bmatrix}$ (5)
Due to the assumption of covariance in $\mathbf{\vec{z}_{\text{var}_{a}}}$ and $\mathbf{\vec{z}_{\text{var}_{b}}}$, it is reasonable to assume that the values $\vec{z}^{\,(x_{\text{low}})}_{\text{var}_{b}}$} and $\vec{z}^{\,(x_{\text{high}})}_{\text{var}_{b}}$} are systematically distinct for distinct values of $\mathbf{\vec{z}_{\text{var}_{a}}}$ under some classes of inputs, $x_{\text{low}}$ and $x_{\text{high}}$, while $\operatorname{sign}({\vec{z}^{\,(x_{\text{low}})}_{\text{var}_{b}}})=\operatorname{sign}({\vec{z}^{\,(x_{\text{high}})}_{\text{var}_{b}}})$. Under these assumptions, if we perform an interchange intervention on $\mathbf{\vec{z}_{\text{var}_{b}}}$ using source representations from input $x_{\text{low}}$ and target representations from input $x_{\text{high}}$, the intervened representation will have values:
$\hat{z}=\begin{bmatrix}\mathbf{\vec{z}_{\text{var}_{a}}}={\vec{z}^{\,(x_{\text{high}})}_{\text{var}_{a}}}\\ \mathbf{\vec{z}_{\text{var}_{b}}}={\vec{z}^{\,(x_{\text{low}})}_{\text{var}_{b}}}\\ \end{bmatrix}$ (6)
Because we assumed that the value of $\vec{z}^{\,(x_{\text{low}})}_{\text{var}_{b}}$ is systematically unique due to covariance in $\mathbf{\vec{z}_{\text{var}_{a}}}$ and $\mathbf{\vec{z}_{\text{var}_{b}}}$, the values $\vec{z}^{\,(x_{\text{high}})}_{\text{var}_{a}}$ and $\vec{z}^{\,(x_{\text{low}})}_{\text{var}_{b}}$ in Equation 6 will have never existed together in the native distribution, but the behavior of the NN will remain unchanged because $\mathbf{\vec{z}_{\text{var}_{b}}}$ is behaviorally binary and its sign has not changed. This is thus a case of harmless divergence.
If we focus only on the divergence in causal subspaces and ignore cases of null-space interactions, what claims can we make about the model’s natural neural mechanisms using divergent, intervened representations? An interpretation consistent with the principles of superposition (Smolensky1988; PDP1; PDP2; elhage2022superposition) is that differences in the exact values of each causal subspace do not matter – rather, only the decision boundaries along these subspaces matter. Under this interpretation, any divergence in the intervened distribution arising from covariance in causal subspaces is acceptable because the separation of the causal subspaces and their decision boundaries (i.e. the computationally important aspects of the NN) are respected in the alignment. This interpretation allows us to ignore computationally irrelevant covariance in our attempt to understand NN mechanisms.
4.2 Pernicious divergence via off-manifold activation
Pernicious cases of representational divergence involve the concept of hidden pathways, which refer to any unit, vector direction, or subcircuit that is inactive on the natural support of representations for a given context but becomes active and influences behavior under an intervention. Although hidden pathways can be compatible with harmless divergences (Section 4.1), they can also undermine claims about natural mechanisms and can prime dormant behavioral changes discussed in Section 4.2.2.
Formally, let $\mathcal{D}$ denote the data distribution over latent representations $h^{\ell}\in\mathbb{R}^{d}$ at layer $\ell$, and let $\mathcal{S}=\mathrm{supp}(h^{\ell}\sim\mathcal{D})$ denote its support. Denoting the intended class $K$ following an intervention as subscripted $\rightarrow K$, an intervened representation $\hat{h}^{\ell}_{\rightarrow K}$ is said to be divergent if $\hat{h}^{\ell}_{\rightarrow K}\notin\mathcal{S}$ or if it causes a ReLU state change for any neuron that does not change sign for the same class K under $\mathcal{S}$. We define the convex hull of class-K representations as $\mathrm{conv}(S_{K})=\{\sum_{i}\alpha_{i}h_{i,K}^{\ell}:\alpha_{i}\geq 0,\,\sum_{i}\alpha_{i}=1\}$ where the subscript $K$ denotes that $h^{\ell}_{i,K}$ was taken from class $K$ inputs. Projecting an intervention onto $\mathrm{conv}(S_{K})$ ensures that it remains within the convex interpolation region of class-K samples.
4.2.1 Mean-difference patching can activate hidden pathways
Claim.
Patching with a mean-difference vector can flip a decision by activating a unit that is silent for all natural class inputs.
Setup.
Consider a two-layer circuit with a ReLU nonlinearity. Let $h^{\ell}\in\mathbb{R}^{4}$ feed into
$s=\mathbf{1}^{\top}h^{\ell+1}=\mathbf{1}^{\top}\mathrm{ReLU}(W_{\ell}h^{\ell}+b_{\ell}),\quad W_{\ell}\in\mathbb{R}^{3\times 4},\,b_{\ell}\in\mathbb{R}^{3},$
where
$W_{\ell}=\begin{bmatrix}0.75&0.25&0&0.5\\ 0&1&0&0\\ 1&1&-1&-1\end{bmatrix},\quad b_{\ell}=\begin{bmatrix}-0.5\\ -0.5\\ 0\end{bmatrix}$.
A positive score ($s>0$) indicates class A. Suppose class-A and class-B representations at layer $\ell$ are
$h^{\ell}_{A}=\begin{cases}[1,0,1,0]^{\top}&\text{(case 1)}\\ [0,1,1,0]^{\top}&\text{(case 2)}\end{cases},\quad h^{\ell}_{B}=\begin{cases}[0,0,1,0]^{\top}&\text{(case 1)}\\ [0,0,1,1]^{\top}&\text{(case 2)}\end{cases}$
Evaluating $h^{\ell}_{A}$ yields
$h^{\ell+1}_{A_{\text{case}_{1}}}=[0.25,0,0]^{\top},\quad h^{\ell+1}_{A_{\text{case}_{2}}}=[0,0.5,0]^{\top},$
so $s_{A}\in\{0.25,0.5\}$. For class B, all outputs are zero, $s_{B}=0$.
Mean-difference patching.
We construct a mean difference vector between the classes,
$\delta_{B\rightarrow A}=\mu_{A}-\mu_{B}=\tfrac{1}{2}\sum_{i=1}^{2}(h^{\ell}_{A_{\text{case}_{i}}})-\tfrac{1}{2}\sum_{i=1}^{2}(h^{\ell}_{B_{\text{case}_{i}}})=[0.5,\,0.5,\,0,\,-0.5]^{\top}.$
Applying this to class B representations gives
$\hat{h}^{\ell}_{B\rightarrow A}=h^{\ell}_{B}+\delta_{B\rightarrow A}=\begin{cases}[0.5,\,0.5,\,1,\,-0.5]^{\top}&\text{(case 1)}\\ [0.5,\,0.5,\,1,\,0.5]^{\top}&\text{(case 2)}\end{cases}$
After propagation through the circuit:
$\hat{h}^{\ell+1}_{B\rightarrow A}=\begin{cases}[0,\,0,\,0.5]^{\top},&\hat{s}_{\text{case}_{1}}=0.5\\ [0.25,\,0,\,0]^{\top},&\hat{s}_{\text{case}_{2}}=0.25\end{cases}$
The intervention flips the decision to class A ($\hat{s}>0$). However, the third hidden unit becomes active only for $\hat{h}^{\ell}_{B\rightarrow A}$, never for natural $h_{A}^{\ell}$. This new activation is a hidden pathway that was silent under all native samples. Thus the mean-difference patch crosses the decision boundary only by activating an off-manifold circuit.
If we project $\hat{h}^{\ell}_{B\rightarrow A}$ onto $\mathrm{conv}(S_{A})$ (or equivalently onto a local PCA subspace of $S_{A}$), this ReLU state change disappears, and the decision boundary is no longer crossed—confirming that the original effect was driven by an off-manifold activation rather than a genuine within-manifold causal mechanism. This experiment does not support claims about the natural mechanisms of the model.
4.2.2 Dormant behavioral changes
Divergent representations can also yield dormant behavioral changes—interventions that appear behaviorally neutral in one context but alter predictions in another. This occurs when an off-manifold activation primes a downstream computation that only manifests under specific contextual inputs.
Illustration.
Extend the network from the previous Section 4.2.1 by adding one row of zeros to $W_{\ell}$ and $b_{\ell}$, producing $h^{\ell+1}\in\mathbb{R}^{4}$ where the final coordinate is always zero. Add a context vector $v\in\mathbb{R}^{4}$ and a final affine layer:
$\hat{y}=W_{\ell+1}(h^{\ell+1}+v)+b_{\ell+1},\quad W_{\ell+1}=\begin{bmatrix}1&1&0.5&0\\ 0&0&0&0\\ 0&0&1&1\end{bmatrix},\;b_{\ell+1}=\begin{bmatrix}0\\ 0.25\\ -1\end{bmatrix}$.
Here, the first argmax index of $\hat{y}=[\hat{y}_{1},\hat{y}_{2},\hat{y}_{3}]^{\top}$ corresponds to class predictions A, B, and C.
Assume $v=[0,0,0,v_{4}]^{\top}$. For $\hat{h}^{\ell+1}_{B_{\text{case}_{1}}\rightarrow A}=[0,0,0.5,0]^{\top}$,
$\hat{y}_{1}=0.25,\quad\hat{y}_{2}=0.25\quad\hat{y}_{3}=(0.5+v_{4})-1=v_{4}-0.5.$
The model predicts class A when $v_{4}<0.75$ but switches to class C when $v_{4}>0.75$. Notably, $v_{4}$ would not naturally cause a class C prediction below a value of $1$ due to the bias threshold. Thus the same intervention that was benign in one context ($v_{4}<0.75$) produces a behavioral flip in another ($1>v_{4}>0.75$) purely due to the latent divergence priming a new pathway.
Implications.
Dormant behavioral changes highlight that behaviorally “safe” interventions can still introduce hidden context dependencies. Detecting them would require evaluating across all possible contexts, which is infeasible in practice. Therefore, causal intervention based experiments should ideally (1) report any introduced representational divergence outside of the null-space and (2) test causal interventions for context-sensitivity.
Summary.
Hidden pathways arise when off-manifold representations activate computations that never occur for natural representations. Such pathways can alter causal conclusions even when behavior appears unchanged depending on the claims. Manifold-preserving projections and ReLU-pattern audits provide practical safeguards against these pernicious forms of divergence.
5 How might we avoid divergent representations?
We have thus far shown that divergent representations are empirically common and that they can be harmless or harmful depending on multiple factors. We now consider the question of how such divergences might be avoided. Some existing methods solve this by constraining counterfactual features (in our case, intervened representations) directly to the natural manifold (verma2024counterfactualexplanationreview). However, we seek a method that generates principled interventions that are constrained to be innocuous. In this pursuit, we first apply the Counterfactual Latent (CL) loss from grant2025mas to the DAS experiments from wu2024alpacadas and find that we can reduce intervened divergence while preserving accuracy on a Llama based LLM. We then introduce a modified CL loss that only targets causal dimensions, and we show that it can improve OOD intervention accuracy.
5.1 Applying the Counterfactual Latent Loss to Boundless DAS
To encourage intervened NN representations to be more similar to the native distribution, we first apply the CL auxiliary loss from grant2025mas to the Boundless DAS setting in wu2024alpacadas. This auxiliary training objective relies on counterfactual latent (CL) vectors as vector objectives. CL vectors are defined as vectors that encode the same causal variable value(s) that would exist in the intervened vector, $\hat{h}$, assuming the interchange intervention was successful. We can obtain CL vectors from sets of natural $h$ vectors from situations and behaviors that are consistent with the values of the CA to which we are aligning. See Figure 3 for a visualization. As an example, assume that we have a CA with causal variables $\text{var}_{u}$, $\text{var}_{w}$, and $\text{var}_{\textit{extra}}$, and following a causal intervention we expect $\hat{h}$ to have a value of $u$ for variable $\text{var}_{u}$ and $w$ for variable $\text{var}_{w}$. A CL vector $h_{\textit{CL}}$ for this example can be obtained by averaging over a set of $m$ natural representations, $h_{\textit{CL}}=\frac{1}{m}\sum_{i=1}^{m}h^{(x_{i})}_{\textit{CL}}$, where each $h^{(x_{i})}_{\textit{CL}}$ has the same variable values: $\text{var}_{u}=u$ and $\text{var}_{w}=w$ (as labeled by the CA).
The CL auxiliary loss $\mathcal{L}_{\textit{CL}}$ introduced in grant2025mas is composed of the mean of an L2 and a cosine distance objective using CL vectors as labels. Using notation defined in Section 2.2, $\mathcal{L}_{\textit{CL}}$ for a single training sample is defined as follows:
$\mathcal{L}_{\textit{CL}}(\hat{h},h_{CL})=\frac{1}{2}\|\hat{h}-h_{\textit{CL}}\|^{2}_{2}-\frac{1}{2}\frac{\hat{h}\cdot h_{\textit{CL}}}{\|\hat{h}\|_{2}\,\|h_{\textit{CL}}\|_{2}}$ (7)
$\mathcal{L}_{\textit{CL}}$ is combined with the DAS behavioral loss $\mathcal{L}_{\textit{DAS}}$ into a single loss term using $\epsilon$ as a tunable hyperparameter: $\mathcal{L}_{total}=\epsilon\mathcal{L}_{\textit{CL}}+\mathcal{L}_{\textit{DAS}}$. The loss is computed as the mean over batches of samples and optimized using gradient descent (Appendix A.4).
Results: We applied the CL loss to the Boundless DAS notebook from wu2024pyvene which reproduces the main result from wu2024alpacadas (see Appendix A.3). Figure 3A provides qualitative visualizations of the decreased divergence when applying the CL loss, and Figure 3B shows both IIA and EMD as a function of increasing the CL weight $\epsilon$. For small values of $\epsilon$, the IIA is maintained (and potentially even improved) while EMD decreases, demonstrating that the CL auxiliary loss can directly reduce divergence in practical interpretability settings without sacrificing the interpretability method.
5.2 Modified CL loss improves OOD interventions in synthetic model settings
We next modify the CL loss to work independently of the DAS behavioral loss and show that it improves OOD intervention performance. We simulate an intermediate layer of an NN by constructing a synthetic dataset of $h$ vectors with known feature dimensions and labels $y$ that we use to train a Multi-Layer Perceptron (MLP). The dataset consists of noisy samples around a set of grid points from two feature dimensions with correlational structure. Specifically, we define a set of features as the Cartesian product of two values along the $x_{1}$-axis $\{-1,1\}$ and five values along the $x_{2}$-axis $\{0,1,2,3,4\}$, resulting in ten unique coordinates that each correspond to one of ten classes. We add noise and covariance to these feature dimensions and concatenate $n$ extra noise dimensions, resulting in simulated vectors $h\in\mathbb{R}^{2+n}$ where $n=16$ unless otherwise stated. The feature dimensions of these vectors are shown as the natural distributions in Figures 1(b) and 3. We then train a small MLP on these representations to predict the class labels using a standard cross entropy loss. After training, we perform DAS analyses with either the behavioral loss or the CL loss independently. See Appendix A.4 for further details on dataset construction, MLP training, and DAS training.
We modify the CL loss by applying it to individual causal subspaces only (as discovered through the DAS training). This allows us to construct $h_{\textit{CL}}^{\text{var}_{i}}$ vectors specific to a single causal variable $\text{var}_{i}$. The modified $\mathcal{L}^{\prime}_{\textit{CL}}$ for a single training sample is defined as follows:
$\hat{h}^{\text{var}_{i}}=\mathcal{A}^{-1}(D_{\text{var}_{i}}\mathcal{A}(\hat{h})),\quad h^{\text{var}_{i}}_{\textit{CL}}=\text{stopgrad}(\mathcal{A}^{-1}(D_{\text{var}_{i}}\mathcal{A}(h_{\textit{CL}})))$ (8)
$\mathcal{L}^{\prime}_{\textit{CL}}=\sum_{i=1}^{n}\mathcal{L}^{\text{var}_{i}}_{\textit{CL}}=\sum_{i=1}^{n}\Big(\frac{1}{2}\|\hat{h}^{\text{var}_{i}}-h^{\text{var}_{i}}_{\textit{CL}}\|^{2}_{2}-\frac{1}{2}\frac{\hat{h}^{\text{var}_{i}}\cdot h^{\text{var}_{i}}_{\textit{CL}}}{\|\hat{h}^{\text{var}_{i}}\|_{2}\,\|h^{\text{var}_{i}}_{\textit{CL}}\|_{2}}\Big)$ (9)
Figure 3: The CL loss reduces representational divergence and can improve out-of-distribution generalization. (a) PCA of natural (orange) and intervened (blue) representations in the Boundless DAS setting presented in wu2024alpacadas for two CL loss weightings with the same final IIA. (b) IIA (orange) and divergence (purple) of intervened representations from Section 5.1 as a function of CL loss weight ($\epsilon$). (c) Diagram of CL loss; rectangles are model representations and $x_{1}$ and $x_{2}$ are deterministic values of the representations along the two synthetic causal dimensions shown in panels (d) and (e). We patch the $x_{2}$ value from source to target using DAS and define the CL representation as the average of all natural representations that possess the same variable values as the post-intervention representation. (d) and (e) two causal feature dimensions of representations from a synthetic dataset consisting of ten classes (colors), with both natural (dark) and intervened (light) representations shown. (d) shows results from DAS trained using behavior only, (e) shows DAS trained using only the CL loss. (f) performance of alignment matrices trained on one task and evaluated on another that uses the same causal dimensions. CL loss leads to higher OOD performance.
Results: Figures 3D and  3E provide a qualitative comparison of intervened and native representations for interventions using a trained DAS rotation matrix. Each dot in the figures shows the values of the feature dimensions for a single representation. The native states are displayed in darker colors and the intervened in lighter. Each hue indicates the ground truth class of the state. We can see a tightening of the intervened representations using the CL loss. Quantitatively, the DAS loss produces EMD values along the feature dimensions of $0.032\pm 0.003$ whereas the CL loss produces $0.007\pm 0.001$ with IIAs of $0.997\pm 0.001$ and $0.9988\pm 0.0005$ respectively on training/test sets with held-out classes.
We hypothesized that reduced divergence could improve IIA when transferring the DAS rotation matrix to OOD settings. To test this, we partitioned the synthetic task into a dense and a sparse cluster of classes based on the relative spacing of 4 distinct classes within each cluster (see Appendix A.4.4 for partition details). We then performed a separate MLP and DAS training on each partition and evaluated the IIA on the held-out partition to produce Figure 3F. The CL loss performs better than the behavioral loss in these OOD settings. We provide CL loss ablations in Appendix A.4.5.
6 Discussion and limitations
In this work we demonstrated that a variety of common causal interventions can produce representations that diverge from a target model’s natural distribution of latent representations. We then showed that although this can have benign effects for many causal claims, it can also activate hidden pathways and trigger dormant behaviors that can perniciously affect other claims. As a step towards mitigating this issue, we provided a broad-stroke solution by directly minimizing the divergence of intervened activity along causal dimensions, mitigating both pernicious and harmless forms of divergence.
A remaining gap in our work is the failure to produce a principled, reliable method for classifying harmful divergence. Additionally, the modified CL loss that we have presented is confined to a narrow set of simplistic settings and is not specific to pernicious divergence. We look forward to exploring ways to classify and mitigate pernicious divergence through self-supervised means in future work.
Where does this leave us with respect to causal interventions in mechanistic interpretability? Given our theoretical findings, any divergence outside of the null-space of NN layers is potentially pernicious. This poses challenges for aspirations of a complete mechanistic understanding of NNs using existing methods alone. However, we note that many practical mechanistic projects can be satisfied by collecting sufficiently large intervention evaluation datasets, and continued development of methods such as the CL loss can reduce the problem even further. We are optimistic for future developments in this field.
7 Acknowledgments
Thank you to the PDP Lab and the Stanford Psychology department for funding. Thank you to Noah Goodman and Jay McClelland for thoughtful feedback and discussions. Thank you to the PDP lab and the Stanford Mech Interp community for opportunities to present and for thoughtful discussion.
8 LLM Usage Statement
We used ChatGPT to elicit editing suggestions for clarity, as well as to improve notational consistency and clarity in the hidden pathways and dormant behavioral changes formalizations in Sections 4.1 and 4.2. We also used ChatGPT to provide an initial layout of the proof offered in Appendix A.2 showing that axis-aligned hyperrectangles are the only manifold shape that do not have divergent source-target vector pairs in coordinate patching settings.
Appendix A Appendix
A.1 Empirical Intervened Divergence Methodological Details
A.1.1 Intervention Methods
We considered three families of interpretability interventions that modify hidden-layer representations. In all three, we visualize the residual stream output from the specified layer:
  1. 1.
    Mean Difference Vector Patching (MDVP) (feng2024languagemodelsbindentities), where an intervention vector $\delta_{\text{MD}}\in\mathbb{R}^{d}$ is defined as the difference in mean activations between two conditions and then added to or subtracted from activations $h\in\mathbb{R}^{d}$. Formally,
    $\hat{h}=h+\delta_{\text{MD}}$ (10)
    We examine the representations $\hat{h}$ from a sample size of 100 unique contexts across 4 token positions at each individual layer. We compare the representations to the native cases of the swapped binding positions. We used layer 10 of Meta’s Meta-Llama-3-8B-Instruct through Huggingface’s transformers package for this task and visualization (touvron2023llama; wolf2019huggingface). We selected layer 10 as it had the lowest EMD difference of all layers, although, we note that this measure did not necessarily correlate with the subjective interpretation of divergence in the qualitative visualizations. We report the EMD difference in Figure 2(a) as the average over all model layers.
  2. 2.
    Sparse Autoencoder (SAE) Projections (bloom2024saelens), where $h$ is projected through a trained encoder $E:\mathbb{R}^{d}\to\mathbb{R}^{k}$ and linear decoder $D:\mathbb{R}^{k}\to\mathbb{R}^{d}$:
    $h^{\prime}=D(E(h)).$ (11)
    SAEs are trained with sparsity penalty $\lambda_{\text{SAE}}$ to encourage interpretable basis functions. We offload further experimental details to the referenced SAElens paper and code base. We compare the reconstructed representations to an equal sample size of 2000 vectors from the natural distribution. We used layer 25 of Meta’s Meta-Llama-3-8B-Instruct through Huggingface’s transformers package for this task and visualization (touvron2023llama; wolf2019huggingface). We selected layer 25 as it appeared to be the only layer available through SAElens’ pretrained SAEs.
  3. 3.
    Distributed Alignment Search (DAS) (wu2024alpacadas), where representations are aligned to a causal abstraction using a learned orthogonal transformation $Q\in\mathbb{R}^{d\times d}$. See Section 2.2 and wu2024alpacadas for further detail on the method. We compare the intervened representations to an equal sample size of 1000 vectors from the natural distribution. We used the model and layer specified in wu2024alpacadas for the visualizations in Figure 2.
A.1.2 Measuring Divergence
For each intervened sample, there exists a corresponding ground truth sample that the intervention is meant to approximate. In the case of the mean difference experiments, these ground samples consist of the naturally occurring entity or attribute in the position which the $\delta_{MD}$ is meant to approximate. For the SAEs, each reconstructed vector has a corresponding encoded vector. For DAS, the ground truth vectors are equivalent to CL vectors.
Earth Mover’s Distance: To quantify distributional differences between original and intervened representations, we approximated the Earth Mover’s Distance (EMD) (villani2009wassersteinEMD) including all vector dimensions using the Sinkhorn loss from the GeomLoss python package with a $p=2$ and $\text{blur}=0.05$ (cuturi2013sinkhorn). Let $\mathcal{H}=\{h_{i}\}_{i=1}^{N}$ denote a set of original representations and $\hat{\mathcal{H}}=\{\hat{h}_{i}\}_{i=1}^{N}$ their intervened counterparts. We computed
$\text{EMD}(\mathcal{H},\hat{\mathcal{H}})=\min_{\gamma\in\Pi(\mu,\nu)}\frac{1}{N}\sum_{i,j}\gamma_{ij}\,\|h_{i}-\hat{h}_{j}\|_{2},$ (12)
where $N$ is the number of samples in $\mathcal{H}$, $\mu$ and $\nu$ are the empirical distributions over $\mathcal{H}$ and $\hat{\mathcal{H}}$, and $\Pi(\mu,\nu)$ denotes the set of couplings with marginals $\mu$ and $\nu$.
Baseline divergence.
To ensure that we did not introduce bias from the sampling procedure in the baseline comparison, we use the corresponding ground truth vectors for the intervened vectors when comparing to the baseline divergence. Formally, let $\mathcal{H}^{\prime}$ be the set of ground truth natural vectors corresponding to the set $\hat{\mathcal{H}}$:
$\text{Divergence}_{\text{Baseline}}=\text{EMD}(\mathcal{H},\mathcal{H}^{\prime}).$ (13)
Local PCA Distance:
For each reference point $x_{i}$, we identified its $k$ nearest neighbors $\mathcal{N}_{k}(x_{i})$ in Euclidean space and computed the local tangent subspace via PCA. Let $U_{d}(x_{i})\in\mathbb{R}^{D\times d}$ denote the top $d$ principal components explaining at least 95% of the local variance. The projection matrix onto this tangent space is $P_{i}=U_{d}U_{d}^{\top}$.
For a query point $v$, the Local PCA Distance is the orthogonal residual between $v$ and its projection onto the local tangent space at the nearest reference sample:
$D_{\mathrm{PCA}}(v)=\min_{x_{i}}\left\|(I-P_{i})\,[\,v-x_{i}\,]\right\|_{2}.$
Small values indicate that $v$ lies close to the locally linear approximation of the manifold, whereas large residuals reflect departures orthogonal to the manifold surface.
Local Linear Reconstruction Error:
We computed an error inspired by Locally Linear Embedding (LLE). Given the same neighborhood $\mathcal{N}_{k}(v)$ of $k$ reference points $X_{k}=\{x_{1},\dots,x_{k}\}$, we found reconstruction weights $w=(w_{1},\dots,w_{k})$ that best express $v$ as a convex combination of its neighbors:
$\min_{w}\ \left\|v-\sum_{j=1}^{k}w_{j}x_{j}\right\|_{2}^{2}\qquad\text{s.t.}\qquad\sum_{j}w_{j}=1.$
A small Tikhonov regularizer $\lambda I$ was added to the local covariance for numerical stability. The Local Linear Reconstruction Error is the residual norm at the optimum:
$D_{\mathrm{LLR}}(v)=\left\|v-X_{k}w^{\ast}\right\|_{2}.$
This metric measures how well $v$ can be expressed as a locally linear interpolation of nearby manifold points; poor reconstruction (large $D_{\mathrm{LLR}}$) implies off-manifold position or local curvature mismatch.
Kernel Density (KDE) Density Score:
We estimated a nonparametric probability density function over the reference set using Gaussian kernel density estimation:
$\hat{p}(x)=\frac{1}{nh^{D}}\sum_{i=1}^{n}\exp\!\left[-\frac{\|x-x_{i}\|_{2}^{2}}{2h^{2}}\right],$
where $h$ is the kernel bandwidth determined by Silverman’s rule of thumb or cross-validation. The KDE Density Score for a query point $v$ is its log-density under this model,
$S_{\mathrm{KDE}}(v)=\log\hat{p}(v),$
which inversely reflects off-manifold distance: lower log-density corresponds to less typical or out-of-distribution points. To express results on a comparable scale, we report the negative log-density (i.e., $S_{\mathrm{KDE}}$), so that larger values consistently indicate greater deviation from the manifold.
A.1.3 Visualization
We visualized both original and intervened representations by projecting onto the first two principal components of the covariance matrix of $\mathcal{H}^{\prime}$ and $\hat{\mathcal{H}}$ combined:
$).\text{PCs}=\text{eigenvectors}\big(\text{Cov}(\begin{bmatrix}\mathcal{H}^{\prime}\\ \hat{\mathcal{H}}\end{bmatrix}$$\big)$. (14)
The top two principal components were used to plot representations in two dimensions, with colors distinguishing intervention method and condition (Figure 2).
Figure 4: A number of additional divergence measures to demonstrate the difference between the natural and intervened distributions. Each is labeled by its y-axis. Each metric is computed over a random sample of natural vectors to simulate the natural manifold, and a sampled set of intervened or natural vectors for which to measure the distance from the natural distribution. We refer to this distribution as the "compared" distribution. The sampled intervened and natural vectors are always the "ground-truth pair" described at the beginning of Appendix A.1. Nearest Cosine Distance: refers to the cosine distance to the nearest sample in the natural manifold. Multiple sampes in the compared distribution can share the same natural sample. This value is averaged over all compared samples. Nearest L2 Distance refers to the cosine distance to the nearest sample in the natural manifold. Multiple sampes in the compared distribution can share the same natural sample. This value is averaged over all compared samples. Min Cos Pairing refers to the lowest cost pairing where cost is the cosine distance between two samples. Vector pairs are exclusive. This value is normalized by the number of samples. Min L2 Pairing refers to the lowest cost pairing where cost is the L2 distance between two samples. Vector pairs are exclusive. This value is normalized by the number of samples. Local PCA Distance refers to the distance to the manifold created using a local PCA of the nearest neighbors. See Appendix A.1.2. EMD refers to the Earth Mover’s Distance. See Appendix A.1.2. KDE refers to the Kernel Density Estimation score. See Appendix A.1.2. Local Linear Reconstruction refers to the local linear reconstruction error. See Appendix A.1.2.
A.2 Closure under coordinate patching characterizes Cartesian products and hyperrectangles
In this section, we will show that axis-aligned hyperrectangles are the only convex manifold shape that does not have source-target vector pairs that produce off-manifold intervened representations.
For vector dimensions $S\subseteq[d]=\{1,\dots,d\}$ and $h^{\textit{src}},h^{\textit{trg}}\in\mathbb{R}^{d}$, define a coordinate patch
$\mathrm{Patch}_{S}(h^{\textit{src}},h^{\textit{trg}})_{i}\;=\;\begin{cases}h_{i}^{\textit{src}},&i\in S,\\ h_{i}^{\textit{trg}},&i\notin S.\end{cases}$
where $i$ refers to the $i^{\text{th}}$ vector coordinate. A set $\mathcal{M}\subseteq\mathbb{R}^{d}$ is patch-closed if $\mathrm{Patch}_{S}(h^{\textit{src}},h^{\textit{trg}})\in\mathcal{M}$ for all $h^{\textit{src}},h^{\textit{trg}}\in\mathcal{M}$ and all $S\subseteq[d]$.
Let $\pi_{i}:\mathbb{R}^{d}\to\mathbb{R}$ be the $i^{\text{th}}$ coordinate projection and write $I_{i}:=\pi_{i}(\mathcal{M})=\{h_{i}:h\in\mathcal{M}\}$.
Theorem A.2 (Patch-closure $\iff$ product of projections).
Let $\mathcal{M}\subseteq\mathbb{R}^{d}$ be nonempty and let $I_{i}:=\pi_{i}(\mathcal{M})$. Then
 $\mathcal{M}\text{ is patch-closed}\iff\mathcal{M}=I_{1}\times\cdots\times I_{d}.$
Proof. ($\Leftarrow$) Immediate: if $\mathcal{M}=\prod_{i}I_{i}$, then patching replaces coordinates by elements of the corresponding $I_{i}$, so the result stays in $\mathcal{M}$.
($\Rightarrow$) Suppose $\mathcal{M}$ is patch-closed. The inclusion $\mathcal{M}\subseteq\prod_{i}I_{i}$ is tautological. For the reverse inclusion, fix $t=(t_{1},\dots,t_{d})$ with $t_{i}\in I_{i}$. For each $i$ pick $h^{(i)}\in\mathcal{M}$ with $h^{(i)}_{i}=t_{i}$. Define $\hat{h}^{(1)}:=h^{(1)}$ and for $k\geq 2$ set $\hat{h}^{(k)}:=\mathrm{Patch}_{\{k\}}(\hat{h}^{(k-1)},h^{(k)})$. Patch-closure gives $\hat{h}^{(k)}\in\mathcal{M}$, and by construction $\hat{h}^{(k)}_{j}=t_{j}$ for all $j\leq k$. Hence $\hat{h}^{(d)}=t\in\mathcal{M}$, yielding $\prod_{i}I_{i}\subseteq\mathcal{M}$. ∎
Corollary (Convex case $\Rightarrow$ hyperrectangle).
If $\mathcal{M}$ is also convex, then each $I_{i}=\pi_{i}(\mathcal{M})\subset\mathbb{R}$ is convex, hence an interval. Therefore $\mathcal{M}=\prod_{i}I_{i}$ is an axis-aligned hyperrectangle (Cartesian product of intervals). Conversely, any axis-aligned hyperrectangle is patch-closed (and convex).
Implication.
Consequently, any nonempty convex set in $\mathbb{R}^{d}$ that is not an axis-aligned hyperrectangle (e.g., a ball, ellipsoid, or a polytope with non-axis-aligned faces) fails to be patch-closed: there exist $h^{\textit{src}},h^{\textit{trg}}\in\mathcal{M}$ and $S\subseteq[d]$ such that $\mathrm{Patch}_{S}(h^{\textit{src}},h^{\textit{trg}})\notin\mathcal{M}$.
A.2.1 Activation patching in balanced subspaces
Here we include an additional example of pernicious activation patching that assumes the existence of balanced subspaces, defined as one or more behaviorally relevant subspaces that are canceled out by opposing weight values. Before continuing, we note that such subspaces are unlikely to exist in practical models due to the fact that they would only arise in cases where two rows of a weight matrix $W\in\mathbb{R}^{n\times m}$ are non-zero, scalar multiples of one another, assuming $h=Wx$. This example, however, could arise in cases where the input $x$ is low rank and a subset of the columns of two rows in $W$ are scalar multiples of one another.
Consider the case where there exists an NN layer that classifies inputs based on the mean intensity of dimensions 3 and 4 for a latent vector $h\in\mathbb{R}^{4}$, where the NN layers that produce $h$ are denoted $f(x)$ with data inputs $x$ sampled from the dataset, $x\sim\mathcal{D}$, and where the layer of interest has a synthetically constructed weight vector $w\in\mathbb{R}^{4}$ where $w^{\top}=\begin{bmatrix}w_{1}&w_{2}&w_{3}&w_{4}\end{bmatrix}=\begin{bmatrix}1&-1&\frac{1}{2}&\frac{1}{2}\end{bmatrix}$. The layer is thus defined as follows:
$y=w^{\top}f(x^{(i)})=w^{\top}h^{(i)}=1h^{(i)}_{1}-1h^{(i)}_{2}+\frac{1}{2}h^{(i)}_{3}+\frac{1}{2}h^{(i)}_{4}$ (15)
[…]
Concretely, if we set $h^{(i)}=[1\;\;1\;\;1\;\;1]^{\top}$ and $h^{(j)}=[3\;\;3\;\;-1\;\;-1]^{\top}$ and then perform an intervention on $h_{2}$ using $h^{(i)}$ as the target and $h^{(j)}$ as the source, we get: $\hat{h}=[1\;\;3\;\;1\;\;1]^{\top}$. This will result in a negative value of $y$, thus crossing its decision boundary using a non-native mechanism. This intervention could be used as experimentally affirming evidence for a mechanistic claim, when in reality we have not addressed the model’s original mechanism.
A.3 CL loss applied to Boundless DAS
To perform the Boundless DAS experiments, we used the Boundless DAS tutorial provided in the pyvene python package (wu2024pyvene) which reproduces a main result from wu2024alpacadas, and we included the CL loss as a weighted auxiliary objective as described in Section 5.1. The exact task used in this tutorial is one involving continuous valued features, which resulted in few occurrences of valid CL vectors in the provided dataset. In order to obtain exact CL vectors, we generated a token sequence sample that contained a valid CL vector for each intervention sample in the dataset. We left hyperparameter choices the same across all trainings except for the CL loss weight $\epsilon$.
A.4 CL Loss in Synthetic Settings
Here we continue Section 5 with experimental details and additional experiments and results.
A.4.1 Synthetic Dataset Construction
The default synthetic task reported in the results section of Section 5.2 was constructed as a dataset of simulated intermediate-layer representations $h\in\mathbb{R}^{18}$ with known ground-truth labels $y\in\{1,\dots,10\}$ and two causal feature dimensions, where 18comes from 2 feature dimensions plus 16 concatenated noise dimensions is the total feature dimensionality. We split these classes into partition 1 and 2, each consisting of 8 of the 10 classes. The held out classes for partition 1 were contained in partition 2 and visa versa. See Figure 5 for a visualization of the task.
Base feature coordinates.
We first defined a grid of base coordinates as the Cartesian product of $\{-1,1\}$ along the first feature axis and $\{0,1,2,3,4\}$ along the second feature axis:
$\mathcal{G}=\{-1,1\}\times\{0,1,2,3,4\},$ (16)
This procedure yields $10=|\mathcal{G}|$ unique base coordinates, each corresponding to a distinct class label.
Figure 5: Visualization of the different synthetic tasks used for Figure 3. The Default Task is split into two partitions, both withholding two classes that are contained in the other partition. The OOD task is also split into two partitions, both consisting of 4 classes. The Dense partition consists of a tighter cluster than the Sparse.
Noise and correlation structure.
For each base coordinate $(x_{1},x_{2})\in\mathcal{G}$, we generated $N$ noisy samples by adding Gaussian noise with variance $0.1^{2}$ and covariance parameter $0.2$. Specifically, each sample was drawn as
[…]
Additional noise dimensions.
We augmented each 2D noisy base vector with $16$ independent Gaussian noise features, each sampled from $\mathcal{N}(0,1)$, producing final representations $h\in\mathbb{R}^{2+16}$.
A.4.2 MLP Training
We trained a feedforward Multi-Layer Perceptron (MLP) classifier to predict the class label $y$ from the synthetic representations $h$. The MLP was parameterized with:
  • input dimensionality $d$, defaulting to $d=18$ as previously described
  • a 1D batch normalization of $d$ dimensions (ioffe2015batchnorm),
  • one hidden layer of width $128$,
  • activation function ReLU,
  • dropout with probability $0.5$ to drop (srivastava14dropout),
  • a 1D batch normalization of $128$ dimensions,
  • output layer with $10$ logits.
Training was performed with a standard categorical cross-entropy loss:
$\mathcal{L}_{\text{CE}}=-\frac{1}{B}\sum_{i=1}^{B}\log p_{\theta}(y_{i}\mid h_{i}),$ (18)
where $B$ is the batch size and $p_{\theta}$ denotes the MLP’s predictive distribution over class labels.
We perform the the MLP training on both partitions combined for the default dataset split. Then we perform an alignment function training on each partition independently and test the alignment on the untrained partition. We report the average IIA over both data partitions for DAS analyses over 5 seeds. Note that we use an independent MLP for each partition in the OOD experiment (Appendix A.4.4.
Optimization used stochastic gradient descent with learning rate $0.01$ for $300$ epochs with early stopping using an Adam optimizer (kingma2017adam). The code was implemented in PyTorch.
A.4.3 DAS Training
Following MLP pretraining, we applied Distributed Alignment Search (DAS) with varying intensities of the behavioral loss and contrastive learning (CL) loss terms. Specifically, the DAS objective was
$\mathcal{L}_{\text{DAS}}=\epsilon_{\text{behavior}}\,\mathcal{L}_{\text{behavior}}+\epsilon_{\text{CL}}\,\mathcal{L}_{\text{CL}},$ (19)
where $\epsilon_{\text{behavior}}$ and $\epsilon_{\text{CL}}$ are tunable coefficients controlling the strength of each term. We only use values of 0 or 1 for $\epsilon_{\text{behavior}}$ and we explore values of $\epsilon_{\text{CL}}$ referring to it as the CL epsilon in figures. We default to an overall learning rate of $0.01$ and subspace size of $1$ unless otherwise specified. Details of these loss functions are provided in Section 5.
Importantly, we stopped training after loss convergence with a patience of 400 training epochs, and we kept the best DAS alignment matrix decided by IIA, and the best CL alignment matrix by EMD. Furthermore, for these trainings, we used a symmetric invertible linear weight matrix as our alignment function as introduced in grant2025das. Namely, the linear alignment matrix $X$ is constructed as $X=(MM^{\top}+\lambda I)S$ where $M\in R^{d_{m}\times d_{m}}$ is a matrix of learned parameters initially sampled from a centered gaussian distribution with a standard deviation of $\frac{1}{d_{m}}$, $I\in R^{d_{m}\times d_{m}}$ is the identity matrix, $\lambda=0.1$ to prevent singular values equal to 0, and $S\in R^{d_{m}\times d_{m}}$ is a diagonal matrix to learn a sign for each column of $X$ using diagonal values $s_{i,i}=\text{Tanh}(a_{i})+\lambda(\text{sign}(\text{Tanh}(a_{i})))$ where each $a_{i}$ is a learned parameter and $\lambda=0.1$ to prevent 0 values. We perform the the alignment function training on both partitions for each synthetic dataset and test each DAS alignment on the untrained partition. We report the average IIA over both data partitions for DAS analyses.
A.4.4 Out-of-Distribution Experimental Details
To perform the OOD CL loss experiments, we partitioned the classes into 2 non-overlapping groups. Two of the 10 classes were excluded entirely. The groups were chosen so that the Sparse set had strictly greater spacing than the Dense set. See a visualization of the Dense and Sparse partitions in Figure 5. A separate MLP training was performed on each partition individually. Then an alignment function was trained on each partition/classifier tuple using the settings specified in Appendix A.4.3. The alignment functions were then tested on the untrained partition. We report IIA values averaged over the performance on each partition.
A.4.5 Further CL Loss Explorations
In these explorations, we explore DAS learning rate and the number of extra noisy dimensions for the OOD experiments. We show accuracies, EMD divergences, and EMD divergences restricted to the causal dimensions. The EMD values are scaled by the number of extra noisy dimensions. We refer to the EMD measurements along causal dimensions only as the Row EMD. We do this for both the trained partitions and held-out partitions for various DAS trainings.
Figure 6: Out of distribution hyperparameter search showing the DAS IIA on OOD validation data for the trained task partition. We see the DAS loss learning rate (lr) and extra concatenated noisy input dimensions (extra_dim) across the panel columns and rows. The DAS+CL reported values include the behavioral loss whereas the CL label excludes the behavioral loss. The pink dashed lines represent DAS trained with the behavioral loss only.
Figure 7: In distribution hyperparameter search showing the DAS IIA on in-distribution validation data for the trained task partition. We see the DAS loss learning rate (lr) and extra concatenated noisy input dimensions (extra_dim) across the panel columns and rows. The DAS+CL reported values include the behavioral loss whereas the CL label excludes the behavioral loss. The pink dashed lines represent DAS trained with the behavioral loss only.
Figure 8: Out of distribution hyperparameter search showing the DAS row-space EMD on OOD validation data for the trained task partition. We see the DAS loss learning rate (lr) and extra concatenated noisy input dimensions (extra_dim) across the panel columns and rows. The DAS+CL reported values include the behavioral loss whereas the CL label excludes the behavioral loss. The pink dashed lines represent DAS trained with the behavioral loss only.
Figure 9: In distribution hyperparameter search showing the DAS row-space EMD on validation data for the trained task partition. We see the DAS loss learning rate (lr) and extra concatenated noisy input dimensions (extra_dim) across the panel columns and rows. The DAS+CL reported values include the behavioral loss whereas the CL label excludes the behavioral loss. The pink dashed lines represent DAS trained with the behavioral loss only.